# Nonseparability of multipartite systems in dilaton black hole

### v1

@@ V Conclusions
The influence of the Hawking effect on N-partite nonseparability of GHZ and W states using the AR $q$-conditional entropy for bosonic and fermionic fields is investigated in the background of the GHS dilaton black hole. Initially, $N$ observers are posited to share GHZ and W states in the asymptotically flat region. The scenario then evolves with $N-1$ observers remaining stationary in the asymptotic region, while a single observer hovers near the event horizon of the black hole. We derive general analytical expressions for the N-partite $q$-conditional entropy of both bosonic and fermionic fields, which quantify the nonseparability of GHZ and W states in the dilaton black hole background. Our findings reveal a striking difference between the behavior of bosonic and fermionic fields under the influence of gravitational effects. Specifically, fermionic nonseparability proves to be more resilient than its bosonic counterpart, despite a weaker fermionic coherence in the dilaton black hole A11 ; A16 . For the bosonic field, both GHZ and W states undergo a transition from nonseparability to separability in the extreme black hole limit. Interestingly, we observe that the nonseparability of GHZ state consistently exceeds that of the W state, while its quantum coherence is weaker in the dilaton black hole. Furthermore, the nonseparability of GHZ state remains unaffected by the number of initial particles $N$, while for W state, the nonseparability exhibits a strong dependence on $N$. As the $N$ increases, the nonseparability of W state decreases, while its quantum coherence grows in the dilaton black hole. A particularly noteworthy result is that the Hawking effect can lead to a net increase in the nonseparability of W state for fermionic fields, a phenomenon that contrasts with the behavior seen in bosonic fields. These results suggest that the choice of quantum resources must be tailored to different particle types and quantum state configurations when handling relativistic quantum information tasks. We hope this research not only enhances our understanding of quantum information behavior in strong gravitational fields but also aids in the effective utilization of quantum resources in future relativistic quantum information applications.

### v2

@@ V Conclusions
The influence of the Hawking effect on N-partite nonseparability of GHZ and W states using the AR $q$-conditional entropy for bosonic and fermionic fields is investigated in the background of the GHS dilaton black hole. Initially, $N$ observers are posited to share GHZ and W states in the asymptotically flat region. The scenario then evolves with N1N-1 observers remaining stationary in the asymptotic region, while a single observer hovers near the event horizon of the black hole. We derive general analytical expressions for the N-partite $q$-conditional entropy of both bosonic and fermionic fields, which quantify the nonseparability of GHZ and W states in the dilaton black hole background. Our findings reveal a striking difference between the behavior of bosonic and fermionic fields under the influence of gravitational effects. Specifically, fermionic nonseparability proves to be more resilient than its bosonic counterpart, despite a weaker fermionic coherence in the dilaton black hole A11 ; A16 . For the bosonic field, both GHZ and W states undergo a transition from nonseparability to separability in the extreme black hole limit. Interestingly, we observe that the nonseparability of GHZ state consistently exceeds that of the W state, while its quantum coherence is weaker in the dilaton black hole. Furthermore, the nonseparability of GHZ state remains unaffected by the number of initial particles $N$, while for W state, the nonseparability exhibits a strong dependence on $N$. As the $N$ increases, the nonseparability of W state decreases, while its quantum coherence grows in the dilaton black hole. A particularly noteworthy result is that the Hawking effect can lead to a net increase in the nonseparability of W state for fermionic fields, a phenomenon that contrasts with the behavior seen in bosonic fields. These results suggest that the choice of quantum resources must be tailored to different particle types and quantum state configurations when handling relativistic quantum information tasks. We hope this research not only enhances our understanding of quantum information behavior in strong gravitational fields but also aids in the effective utilization of quantum resources in future relativistic quantum information applications.